Please answer these question:--- Calculate the Ka2 for oxalic acid (found in rhubarb and spinach) (H2C204) if a 0.35 M solution of Na2C204 has a pH of 8.8
Answer #1

part 1

Na2C2O4 is salt of strong base ( NaOH) and weak acid (NaHC2O4) .

given concentration of salt (Csalt) = 0.35 M

hence,

pH = 7 + \frac{1}{2} ( pKa2 + logCsalt )

pH = 7 + \frac{1}{2} ( pKa2 + log 0.35)

or, 8.87 = 7 + \frac{1}{2} ( pKa2 - 0.456)

or, pKa2 - 0.456 = 2 × (8.87 - 7)

or, pKa2 - 0.456 = 3.74

or, pKa2 = 3.74 + 0.456 = 4.196.

hence Ka2 of oxalic acid  = 10-4.196 = 6.36×10-5

so,

part 2

given , Kb = 7.4×10-7 , hence,

pKb = - logKb = - log (7.4×10-7) = 6.13

pH = 10.72

for weak base

pH =14 - \frac{1}{2} ( pKb - logC)

or, 10.72 = 14 - \frac{1}{2} ( 6.13 - log C)

or, -\frac{1}{2} (6.13 - logC) = (10.72 - 14)

or, - (6.13 - log C) = - (3.28×2)

or, log C - 6.13 = - 6.56

or, log C = (6.13 - 6.56)

or, log C = - 0.43

or, C = 10(-0.43)

or, C = 0.37 M.

concentration of nicotine = 0.37 M.